# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
and by definition of $E_{w_{i,k}}(R)$, we have
$g=$ $(a_{-t_{k}}f_{i,k}^{-1}a_{t_{k}})\cdot w_{i,k}^{-1}\cdot q_{k}$
$\in$ $\left(a_{-t_{k}}\cdot E_{w_{i,k}}\left(C_{1}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w_{i,k}^{-1}\cdot q_{k}$
for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho$, we know that
$\|\rho_{\beta_{0}}(w_{i,k}f_{i,k}^{-1}w_{i,k}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w_{i,k}}(f_{i,k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
@@
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
@@
$d(q_{k})\leq C_{2}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i,k}^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By passing to a subsequence, we may assume that $w_{i,k}$ is a fixed element in $\{w_{i}\}_{i\in\bar{I}}$ independent of $k$. In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
@@
For any $i\in\bar{I}$, define $\mathcal{E}_{i}(\tau)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
@@
Let $i\in\bar{I}$. Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $\theta_{i}>0$ depending only on $i$ such that $d(q)\geq\theta_{i}$.
@@
By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$ By Corollary 2.3, let
$w_{i}^{-1}\cdot q=u\cdot v$
for some $u\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Let $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ be a bounded fundamental domain of $\mathbf{F}_{w_{i}}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R}))$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$. We may write
@@
for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w_{i}}(\mathbb{R})$. Then
$q\cdot\mathbb{Z}^{d}=(w_{i}\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(w_{i}uw_{i}^{-1})\cdot(w_{i}\tilde{v}\cdot\mathbb{Z}^{d}).$
By the fact that $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, we can conclude that the co-volume of $(w_{i}\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w_{i}$ and $\tilde{\Omega}_{\mathbf{F}_{w_{i}}}$, and so does $q\cdot\mathbb{Z}^{d}$. This completes the proof of the lemma. ∎
Now we choose $\epsilon_{0}>0$ so that
$\min_{i\in\bar{I}}\{\theta_{i}\}>C_{0}\cdot\epsilon_{0}^{\dim V_{\beta_{0}}}.$
@@
If $\mathcal{E}_{i}(\tau)\neq\emptyset$ for some $i\in\bar{I}$, then $0\leq\tau<\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$.
@@
Let $g\in\mathcal{E}_{i}(\tau)$. By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}$
$d(q_{k})\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
$\theta_{i}\leq C_{0}(\epsilon_{0}\cdot e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
If $\tau\geq\beta_{0}(w_{i}a_{-1}w_{i}^{-1})$, then by taking $t_{k}\to\infty$, we obtain that
$\theta_{i}\leq C_{0}\cdot(\epsilon_{0})^{\dim V_{\beta_{0}}}$
which contradicts the choice of $\epsilon_{0}$. This completes the proof of the lemma. ∎
From the discussion above, we obtain the following
@@
We have
$S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{i\in\bar{I}}\mathcal{E}_{i}(\tau).$
In the following, for each $i\in\bar{I}$, we will consider the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ $(\mathcal{E}_{i}(\tau)\neq\emptyset)$ and compute an upper bound for the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. These upper bounds will give an upper bound for $\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6. For notational convenience, we may set $\epsilon_{0}=1$ as it does not affect the computations in the rest of this section.
Now fix $i\in\bar{I}$. To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we first need to estimate the volume of the subset $E_{w_{i}}(R)$ in ${\mathbf{F}}_{w_{i}}(\mathbb{R})$, which can be written as
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))$
where $B_{R}$ is the ball of radius $R$ centered at 00 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w_{i}}}$ is the Haar measure on ${\mathbf{F}}_{w_{i}}(\mathbb{R})$. We need the following result about the asymptotic volume estimates of algebraic varieties.
Theorem 6.7 ([3, Corollary 16.3]).
[…]
@@
We apply Theorem 6.7 to the morphism $\Psi_{w_{i}}$ where the closed orbit $\mathcal{O}=\Psi_{w_{i}}(\mathbf{F}_{w_{i}}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w_{i}}^{*}(\mu_{\mathbf{F}_{w_{i}}})$ of the Haar measure on $\mathbf{F}_{w_{i}}(\mathbb{R})$, and obtain the following
@@
There exist constants $a_{w_{i}}\in\mathbb{Q}_{\geq 0}$, $b_{w_{i}}\in\mathbb{Z}_{\geq 0}$ and $c_{w_{i}}>0$ such that
$\operatorname{Vol}(E_{w_{i}}(R))=\mu_{\mathbf{F}_{w_{i}}}(\Psi_{w_{i}}^{-1}(B_{R}))\sim c_{w_{i}}R^{a_{w_{i}}}(\log R)^{b_{w_{i}}}\;(\textup{as }R\to\infty).$
To compute the Hausdorff dimension of $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$, we also need an upper bound for the asymptotic number of rational elements in
$(w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
@@
$\tilde{\Psi}_{w_{i}}:w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{R})\to\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V,\quad\tilde{\Psi}_{w_{i}}(x)=\rho_{\beta_{0}}(x)\cdot e_{V_{\beta_{0}}}.$
@@
Let $g\in w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in a bounded set $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$. By definition, $d(g)$ is the co-volume of $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C>0$ depending only on $\tilde{U}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C\cdot d(g)$. So the number of rational points $g$ in $\tilde{U}$ whose denominators are less than $l>0$ is less than or equal to the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C\cdot l$. This leads us to the results about the Manin’s conjecture [2] (for relevant results about the Manin’s conjecture, one may refer to [11, 16, 18, 19, 32, 36, 38, 40] and the references therein.) In particular, there exists a constant $A_{w_{i}}>0$ such that for any $\epsilon>0$
($\ast$) #|$\#\left|\{g\in\tilde{U}\cap w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}(\mathbb{Q}):d(g)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{A_{w_{i}}+\epsilon}$

### v2

@@ 6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
$\|\rho_{\beta_{0}}(wf_{k}^{-1}w^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim$ $\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|\lesssim C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k})$
and by definition of $E_{w}(R)$, we have
$g=$ $(a_{-t_{k}}f_{k}^{-1}a_{t_{k}})\cdot w^{-1}\cdot q_{k}$
$\in$ $\left(a_{-t_{k}}\cdot E_{w}\left(C_{1}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k})\right)\cdot a_{t_{k}}\right)\cdot w^{-1}\cdot q_{k}$
for some constant $C_{1}>0$ depending only on $\mathbf{G}$, $\Gamma$ and $\rho$. Moreover, by the structure of the representation $\rho_{\beta_{0}}$, we know that
$\|\rho_{\beta_{0}}(wf_{k}^{-1}w^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\Psi_{w}(f_{k}^{-1})\|\gtrsim\|e_{V_{\beta_{0}}}\|=1$
@@
$0pt(q_{k})\leq C_{2}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
@@
$0pt(q_{k})\leq C_{2}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
and $w^{-1}\cdot q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. In the following, we take $C_{0}=\max\{C_{1},C_{2}\}$.
@@
For any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and any $\epsilon>0$, define $\mathcal{E}_{w,\epsilon}(\psi)$ to be the subset of elements $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w}(C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)w^{-1}q_{k}$
$0pt(q_{k})\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
and $w^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
@@
Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ with $w^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $\theta_{w}>0$ depending only on $w$ such that $d(q)\geq\theta_{w}$.
@@
By definition, we know that $R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$ By Corollary 2.3, let
$w^{-1}\cdot q=u\cdot v$
for some $u\in\mathbf{H}_{w}(\mathbb{R})$ and $v\in\mathbf{F}_{w}(\mathbb{Q})$. Let $\tilde{\Omega}_{\mathbf{F}_{w}}$ be a bounded fundamental domain of $\mathbf{F}_{w}(\mathbb{R})/(\Gamma\cap\mathbf{F}_{w}(\mathbb{R}))$ in $\mathbf{F}_{w}(\mathbb{R})$. We may write
@@
for some $\tilde{v}\in\tilde{\Omega}_{\mathbf{F}_{w}}$ and $\gamma\in\Gamma\cap\mathbf{F}_{w}(\mathbb{R})$. Then
$q\cdot\mathbb{Z}^{d}=(w\cdot u\cdot v)\cdot\mathbb{Z}^{d}=(wuw^{-1})\cdot(w\tilde{v}\cdot\mathbb{Z}^{d}).$
By the fact that $w\mathbf{H}_{w}w^{-1}$ fixes every element in $V_{\beta_{0}}$ and $\tilde{v}$ is in the bounded subset $\tilde{\Omega}_{\mathbf{F}_{w}}$, we can conclude that the co-volume of $(w\tilde{v}\cdot\mathbb{Z}^{d})\cap V_{\beta_{0}}$ has a lower bound depending only on $w$ and $\tilde{\Omega}_{\mathbf{F}_{w}}$, and so does $q\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$. This completes the proof of the lemma. ∎
@@
Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $\epsilon>0$. If $\mathcal{E}_{w,\epsilon}(\psi)\neq\emptyset$, then $\tau-\epsilon\leq\beta_{0}(wa_{-1}w^{-1})$.
@@
Let $g\in\mathcal{E}_{w,\epsilon}(\psi)$. By definition, there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g\in\left(a_{-t_{k}}E_{w}(C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)w^{-1}q_{k}$
$0pt(q_{k})\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
and $w^{-1}q_{k}\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 6.4, we have
$\theta_{w}\leq C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}\cdot e^{\lambda_{w}(\epsilon)\cdot t_{k}}.$
Then by taking $t_{k}\to\infty$, we obtain that $\tau-\epsilon\leq\beta_{0}(wa_{-1}w^{-1})$. This completes the proof of the lemma. ∎
@@
For any $\epsilon>0$, we have
$S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathcal{E}_{w,\epsilon}(\psi).$
Moreover, for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, $\beta_{0}(wa_{-1}w^{-1})\leq\beta_{0}(a_{-1})$.
Proof.
The inclusion
$S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\subseteq\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathcal{E}_{w,\epsilon}(\psi)$
follows from the discussion above. Note that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, $\beta_{w}(a):=\beta_{0}(waw^{-1})$ $(\forall a\in\mathbf{T})$ is a weight in the representation $\rho$, and $\beta_{0}$ is the highest weight. It implies that $\beta_{w}(a_{-1})\leq\beta_{0}(a_{-1})$. This completes the proof of the proposition. ∎
In the following, we will consider the subsets $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ $(w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}})$ and compute upper bounds for the Hausdorff dimensions of these subsets. These upper bounds will give an upper bound for $\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ by Proposition 6.6. Note that we are only interested in the non-empty subsets $\mathcal{E}_{w,\epsilon}(\psi)\neq\emptyset$. So here we will only consider the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ which satisfy $\tau\leq\beta_{0}(wa_{-1}w^{-1}).$ For the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ with $\tau>\beta_{0}(wa_{-1}w^{-1})$, we may choose $\epsilon>0$ sufficiently small so that $\tau-\epsilon>\beta_{0}(wa_{-1}w^{-1})$, and then by Lemma 6.5, $\mathcal{E}_{w,\epsilon}(\psi)=\emptyset$. For notational convenience, we may set $C_{0}\cdot C_{\epsilon}^{\dim V_{\beta_{0}}}=1$ as it does not affect the computations in the rest of this section.
Now let $\epsilon>0$ be a sufficiently small number and fix $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ with $\tau\leq\beta_{0}(wa_{-1}w^{-1})$. To compute the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$, we first need to estimate the volume of the subset $E_{w}(R)$ in ${\mathbf{F}}_{w}(\mathbb{R})$, which can be written as
$\operatorname{Vol}(E_{w}(R))=\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))$
where $B_{R}$ is the ball of radius $R$ centered at 0 in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ and $\mu_{\mathbf{F}_{w}}$ is the Haar measure on ${\mathbf{F}}_{w}(\mathbb{R})$. We need the following result about the asymptotic volume estimates of algebraic varieties.
Theorem 6.7 ([4, Corollary 16.3]).
Let $\mathcal{O}$ be a closed orbit of a group $\mathbf{H}(\mathbb{R})$ of real points of an algebraic group $\mathbf{H}$ in an $\mathbb{R}$-vector space $V$, $\mu$ an $\mathbf{H}(\mathbb{R})$-invariant measure on $\mathcal{O}$ and $\|\cdot\|$ a Euclidean norm on $V$. Let $B_{R}=\{v\in V:\|v\|\leq R\}$. Then
@@
We apply Theorem 6.7 to the morphism $\Psi_{w}$ where the closed orbit $\mathcal{O}=\Psi_{w}(\mathbf{F}_{w}(\mathbb{R}))$ and the measuer $\mu$ is the push-forward $\Psi_{w}^{*}(\mu_{\mathbf{F}_{w}})$ of the Haar measure on $\mathbf{F}_{w}(\mathbb{R})$, and obtain the following
@@
There exist constants $a_{w}\in\mathbb{Q}_{\geq 0}$, $b_{w}\in\mathbb{Z}_{\geq 0}$ and $c_{w}>0$ such that
$\operatorname{Vol}(E_{w}(R))=\mu_{\mathbf{F}_{w}}(\Psi_{w}^{-1}(B_{R}))\sim c_{w}R^{a_{w}}(\log R)^{b_{w}}\;(\textup{as }R\to\infty).$
To compute the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$, we also need an upper bound for the asymptotic estimate of the number of rational elements in
$(w\mathbf{F}_{w}w^{-1})(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q}).$
@@
$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
@@
Let $\Omega_{\beta_{0}}$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ which contains a fundamental domain of $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\Gamma)$. Let $g\in w\mathbf{F}_{w}w^{-1}(\mathbb{Q})\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\Omega_{\beta_{0}}$. By definition, $0pt(g)$ is the co-volume of $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, this implies that there exists a constant $C_{\beta_{0}}>0$ depending only on $\Omega_{\beta_{0}}$, such that the length of the primitive integral vector in the line spanned by $\rho(g^{-1})\cdot e_{V_{\beta_{0}}}$ is less than $C_{\beta_{0}}\cdot 0pt(g)$. So the number of rational elements $g$ in $\Omega_{\beta_{0}}\cap(w\mathbf{F}_{w}w^{-1})(\mathbb{Q})$ whose heights are less than $l>0$ is bounded above by the total number of primitive integral points in $\bigwedge_{i=1}^{\dim V_{\beta_{0}}}V$ whose lengths are less than $C_{\beta_{0}}\cdot l$. This leads to the results about Manin’s conjecture [3], i.e. the asymptotic estimate of the number of rational points in an algebraic variety (See e.g. [12, 24, 33, 34, 63, 68, 72, 76, 73] and the references therein.) For the general situation, we define
$A_{w}=A_{w}(\Omega_{\beta_{0}}):=\limsup_{l\to\infty}\frac{\log|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|}{\log l}$
and from the discussion above, we know that $A_{w}$ is a finite number which is bounded above by the growth rate of the number of the primitive integral points in $\bigwedge^{\dim V_{\beta_{0}}}V$. We also have that for any $\epsilon>0$
$|\{g\in\Omega_{\beta_{0}}\cap w\mathbf{F}_{w}w^{-1}(\mathbb{Q}):0pt(g)\leq l\}|\lesssim_{\epsilon,\Omega_{\beta_{0}}}l^{A_{w}+\epsilon}$